Suppose that you decide to hold the bonds in the previous question for a year, and then sell them. The next year, the interest rate on comparable bonds increases to 9.5%. How much will you be able to sell them for at that time? If you received a 5% rate of return on your money and you stick the coupons in the bank, in retrospect, would you have been better off selling the bonds when you first got them, or, holding them for a year, and selling them at the higher yield (9.5%)?  

Prev Question: 2.It is your lucky day: you have been fired from your job. But when your stupid boss kicked you out of the office, he accidentally handed you two $100,000-dollar bearer bonds (non-registered bonds – you only need to present the coupons for payment) with the stack of papers that he cleared from your desk. The bonds have a 10% annual coupon and maturity of five years. If the current market rate on the bonds is 8.7%, what is this severance package worth?  
Answer #1

Value of the bond can be calculated using PV function on a calculator

N = 5, I/Y = 8.7%, PMT = 10% x 100,000 = 10,000, FV = 100,000

=> Compute PV = $105,096.15 is the price of a the bond a year earlier.

Today, I/Y = 9.5%, N = 4 => Compute PV = $101,602.24 is the bond price today.

But you would also receive $10,000 in coupon, and hence the total value of the bond today = $111,602.24

Hence, your returns = 111,602.24 / 105,096.15 - 1 = 6.19% > 5%

As the bond price declined in the last one year but your total returns are still higher than 5% and hence, you are better to hold them for a year.